OpenStudy (anonymous):

What is the equation of a line that passes through the point (-2,1) and (1,10)? 3X-Y=-5 X+3Y=-5 3X-Y=-7 3X-Y=5

OpenStudy (shamim):

required equation is y-y1=((y2-y1)/(x2-x1))(x-x1)

OpenStudy (shamim):

x1=-2 y1=1

OpenStudy (anonymous):

\[\frac{ y2-y1 }{ x2-x1 }=\frac{ y-y1 }{ x-x1 }\]

OpenStudy (shamim):

x2=1 y2=10

OpenStudy (anonymous):

can you please keep going so I can see this problem done as an example? Thanks!

OpenStudy (shamim):

slope of ur line is m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

I dont understand?

OpenStudy (shamim):

y-1=(9/3)(x+2) y-1=3(x+2) y-1=3x+6 y=3x+7

OpenStudy (shamim):

3x-y=-7

OpenStudy (shamim):

c my work once again